# UAV-DETR: Efficient End-to-End Object Detection for Unmanned Aerial Vehicle Imagery

### v1

@@ Abstract
Unmanned aerial vehicle object detection (UAV-OD) has been widely used in various scenarios. However, most existing UAV-OD algorithms rely on manually designed components, which require extensive tuning. End-to-end models that do not depend on such manually designed components are mainly designed for natural images, which are less effective for UAV imagery. To address such challenges, this paper proposes an efficient detection transformer (DETR) framework tailored for UAV imagery, i.e., UAV-DETR. The framework includes a multi-scale feature fusion with frequency enhancement module, which captures both spatial and frequency information at different scales. In addition, a frequency-focused down-sampling module is presented to retain critical spatial details during down-sampling. A semantic alignment and calibration module is developed to align and fuse features from different fusion paths. Experimental results demonstrate the effectiveness and generalization of our approach across various UAV imagery datasets. On the VisDrone dataset, our method improves AP by 3.1% and $\text{AP}_{50}$ by 4.2% over the baseline. Similar enhancements are observed on the UAVVaste dataset. The project page: https://github.com/ValiantDiligent/UAV-DETR
@@ I Introduction
Figure 1: Challenges in UAV-OD
To handle the challenge of object detection in aerial imagery, this paper proposes an efficient detection transformer framework for UAV imagery, namely UAV-DETR. We enhance our model by leveraging both spatial and frequency domain information across multiple scales to retain high-frequency components. We present a frequency-focused down-sampling strategy to preserve critical spatial details during down-sampling. Finally, we enhance the semantic representation capability of the model by aligning features from different feature fusion paths.
@@
    We propose UAV-DETR, an efficient end-to-end detector transformer for UAV imagery. The framework achieves superior accuracy and real-time performance.
@@
    We develop a frequency-focused down-sampling module that retains dual-domain information.
@@ III Methodology
As shown in Fig. 2, this study proposes a UAV-DETR, which is built upon the architecture of RT-DETR [6]. We enhance the model with three components, i.e., multi-scale feature fusion with frequency enhancement, frequency-focused down-sampling, and semantic alignment and calibration. Additionally, we introduce inner Scylla intersection over union (Inner-SIoU) [18, 19] to replace the generalized intersection over union (GIoU).
Figure 2: Overview of the UAV-DETR. MSFF-FE represents the multi-scale feature fusion with frequency enhancement module; FD denotes frequency-focused down-sampling; SAC is semantic alignment and calibration. FFT and IFFT denote fast Fourier transform and its inverse operation, respectively.
@@ III-A Multi-Scale Feature Fusion with Frequency Enhancement
Figure 3: Structure of multi-scale feature fusion with frequency enhancement.
As depicted in Fig. 3, the MSFF-FE adopts a cross stage partial strategy [20]. It partitions the input feature map $x\in\mathbb{R}^{C\times H\times W}$ into two parts: $x_{1}\in\mathbb{R}^{C_{1}\times H\times W}$ and $x_{2}\in\mathbb{R}^{C_{2}\times H\times W}$, where $C_{1}=\frac{C}{4}$ and $C_{2}=\frac{3C}{4}$. The first part $x_{1}$ undergoes multi-scale and frequency enhancement, while the second part $x_{2}$ is concatenated with the processed features from $x_{1}$, followed by a 1×1 convolution to fuse the two branches. Firstly, the feature map $x_{1}$ passes through a 1×1 convolution to adjust its channel dimensions, followed by the GELU activation function[21] for non-linearity, resulting in $x_{\text{conv}}$. Then, frequency domain enhancement is applied using a global average pooling (GAP) operation, followed by a Fourier Transform and an inverse Fourier Transform.
@@
The multi-scale features $x_{\text{sc}}$ are further refined by applying a channel attention mechanism. To enhance frequency-based details, we leverage the gating mechanism to modulate and refine the multi-scale features.
@@
where $\alpha$ and $\beta$ are learned parameters balancing the spatial and frequency components. We refer to this equation as the frequency-focused module, which will be used in later stages of the network. Before combining with the unprocessed $x_{2}$, the enhanced features undergo final fusion.
@@
The final output is obtained by concatenating $x_{\text{final}}$ with $x_{2}$, followed by a 1×1 convolution.
III-B Frequency-Focused Down-Sampling
As illustrated in Fig. 2, in the frequency-focused down-sampling module (FD), the input feature map $x\in\mathbb{R}^{C\times H\times W}$ is first down-sampled using average pooling with a kernel size of 2 and a stride of 1. As a result, the pooled feature map $x_{p}$ is obtained and then it is divided into two parts: $x_{1}$ and $x_{2}$. Each of them are processed in parallel.
$x_{1}$ is processed using a 3×3 convolution with stride 2 and padding 1, which reduces its spatial dimensions while preserving key features. Then, $x_{1}^{\prime}$ is obtained. $x_{2}$ undergoes parallel processing, where one path applies frequency-focused module to enhance important feature components. Accordingly, $x_{f}$ is obtained. The other path applies max pooling with a 3×3 kernel and stride 2, followed by a 1×1 convolution to reduce the number of channels. Correspondingly, we get $x_{p}^{\prime}$. The two outputs, $x_{f}$ and $x_{p}^{\prime}$, are concatenated along the channel dimension and passed through a 1×1 convolution to reduce the number of channels to the desired size, which results in $x_{2}^{\prime}$. Finally, $x_{1}^{\prime}$ and $x_{2}^{\prime}$ are concatenated to form the final output of the module.
Figure 4: Structure of semantic alignment and calibration.
@@ III-C Semantic Alignment and Calibration
Given two input features $x_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ and $x_{2}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$, the SAC module first unifies the number of channels to a common dimension $C$ through separate convolution layers. Then, the feature $x_{2}$ is upsampled using bilinear interpolation to match the spatial dimensions of $x_{1}$. To enhance feature $x_{2}$, we apply the frequency-focused module, which selectively amplifies high-frequency components. It generates a frequency-enhanced feature $x_{\text{freq}}$. Afterward, we fuse the frequency-enhanced feature with the original upsampled feature $x_{2}$. A gating mechanism is employed to balance the contributions from both the spatial and frequency domains.
@@
To address misalignment between $x_{1}$ and $x_{\text{fused}}$, the SAC module learns 2D offsets $\Delta_{1}$ and $\Delta_{2}$, which adjust the sampling grid for each feature map. These offsets are generated through a convolutional layer. Using the learned offsets, we adjust the spatial coordinates of the features through a grid-based sampling operation [22] to ensure the alignment of both features.
@@
Then, an element-wise weighted summation is used to fuse the aligned features, where $\alpha_{1}$ and $\alpha_{2}$ are learned attention weights that balance the contributions from each aligned feature.
$x_{\text{output}}=\alpha_{1}\cdot x_{1}^{\text{aligned}}+\alpha_{2}\cdot x_{\text{fused}}^{\text{aligned}}$ (8)
@@ III-D LOSS Function
RT-DETR uses GIoU loss for bounding box regression, which is less effective for small object detection, especially when the Intersection over Union (IoU) values are low. To address this issue, we adopt Inner-SIoU, which enhances both small object detection and geometric alignment. Inner-SIoU combines Inner-IoU[19] and SCYLLA-IoU (SIoU)[18]. It scales the auxiliary bounding box by 1.25 to improve sensitivity and speed up convergence. Like SIoU, it adds angle and shape losses to reduce angular and distance mismatches.
The InnerIoU is defined using expanded auxiliary bounding boxes. For a given predicted box $B^{\text{inner}}$ and ground truth box $B^{\text{inner}}_{\text{gt}}$, the Inner-IoU is calculated as:
@@ IV-A Experimental Setup
Datasets: We conduct quantitative experiments on two object detection datasets: VisDrone[23] and UAVVaste[24]. The VisDrone-2019-DET dataset comprises 6,471 training images, 548 validation images, and 3,190 test images, all captured from drones at varying altitudes across different locations. Each image is annotated with bounding boxes for ten predefined object categories: pedestrian, person, car, van, bus, truck, motorbike, bicycle, awning-tricycle, and tricycle. We used the VisDrone-2019-DET training set and validation set for training and testing, respectively.
@@
Implementation Details: All models are trained on NVIDIA GeForce RTX 3090. For the details of the network architecture, our UAV-DETR model is based on RT-DETR[6], with two model sizes designed: one using ResNet18 and the other using ResNet50 as the backbone. We implement our approach and train the framework for 400 epochs with a batch size of 4. We use an early stopping mechanism with a patience setting of 20. The UAV-DETR network is optimized using AdamW[25] with a learning rate of 0.0001, a momentum of 0.9. We scale our input images to 640$\times$640 pixels and use the data augmentation methods from the RT-DETR model to ensure consistency across experiments. Additionally, we apply mixup [26] and Mosaic [27] techniques, with Mosaic set to a probability of 1 and mixup set to a probability of 0.2. We report the standard COCO metrics, including AP (averaged over uniformly sampled IoU thresholds ranging from 0.50-0.95 with a step size of 0.05), and AP50 (AP at an IoU threshold of 0.50).
IV-B Comparative Experiments
As listed in Table I, on the VisDrone dataset, our UAV-DETR-R18 achieves a 3.1% improvement in AP, a 4.2% increase in $\text{AP}_{50}$ compared to the baseline RT-DETR-R18. Similarly, the UAV-DETR-R50 sees a 3.1% increase in AP, a 4.1% rise in $\text{AP}_{50}$ compared to the baseline. UAV-DETR-R18 outperforms all methods with a computational cost below 100 GFLOPs, achieving the best accuracy in its class. Furthermore, we compared our method with other object detectors that have similar computational costs to UAV-DETR-R50, and the results show that our approach also outperforms the others in terms of accuracy. Remarkably, our method still exhibits outstanding performance, even compared to approaches like PP-YOLOE-P2-Alpha-l [28], which typically benefit from extensive pre-training. Such comparisons are often considered unfair due to the significant advantages conferred by extensive pre-training.
@@ IV-B Comparative Experiments
YOLOv9-M[30] arXiv2024 640×640 20.1 76.8 25.2 42.0
@@
DetectoRS w/ RFLA[33] ECCV2022 800×800 123.2 160 27.4 45.3
HIC-YOLOv5[2] ICRA2024 640×640 9.4 31.2 20.8 36.1
@@
Model AP AP50
HIC-YOLOv5[2] 40.5 66.8
YOLOv8-M[29] 40.5 68.6
RT-DETR[6] 42.4 71.1
UAV-DETR(Ours) 45.7 74.7
To further demonstrate the generalization ability of UAV-DETR, we also evaluated the method on the UAVVaste dataset. UAV-DETR-R18 was chosen for this comparison as it strikes an optimal balance between computational efficiency and detection accuracy, making it well-suited for evaluation on a smaller dataset like UAVVaste, where maintaining strong performance with limited data is crucial. The compared methods were implemented using the ultralytics toolkits[29], and the results are shown in Table II. Notably, UAV-DETR still maintains a competitive advantage compared to other models, with an improvement of 3.3% and 3.6% in AP and $\text{AP}_{50}$, respectively, compared to the baseline. The results indicate that the proposed method is both feasible and effective for detecting objects in drone images. Compared to VisDrone, the UAVVaste dataset has a smaller amount of data. The strong performance of our model demonstrates that it does not rely on large amounts of annotated data.
@@
IV-C Ablation Studies
We conducted ablation experiments on the VisDrone dataset using UAV-DETR-R18 to analyze the impact of each component on detection accuracy. Table III presents the performance comparison with different configurations, where IS stands for Inner-SIoU, MSFF-FE represents Multi-Scale Feature Fusion with Frequency Enhancement module, FD stands for Frequency-Focused Down-Sampling module, and SAC refers to Semantic Alignment and Calibration module.
@@ IV-C Ablation Studies
RT-DETR-R18[6] 20 60 85
RT-DETR-R50[6] 42 130 40
UAV-DETR-R18 20 77 51
UAV-DETR-R50 42 170 30

### v3

@@ Abstract
Unmanned aerial vehicle object detection (UAV-OD) has been widely used in various scenarios. However, most existing UAV-OD algorithms rely on manually designed components, which require extensive tuning. End-to-end models that do not depend on such manually designed components are mainly designed for natural images, which are less effective for UAV imagery. To address such challenges, this paper proposes an efficient detection transformer (DETR) framework tailored for UAV imagery, i.e., UAV-DETR. The framework includes a multi-scale feature fusion with frequency enhancement module, which captures both spatial and frequency information at different scales. In addition, a frequency-focused downsampling module is presented to retain critical spatial details during downsampling. A semantic alignment and calibration module is developed to align and fuse features from different fusion paths. Experimental results demonstrate the effectiveness and generalization of our approach across various UAV imagery datasets. On the VisDrone dataset, our method improves AP by 3.1% and $\text{AP}_{50}$ by 4.2% over the baseline. Similar enhancements are observed on the UAVVaste dataset. The project page: https://github.com/ValiantDiligent/UAV-DETR
@@ I Introduction
To handle the challenge of object detection in aerial imagery, this paper proposes an efficient detection transformer framework for UAV imagery, namely UAV-DETR. We enhance our model by leveraging both spatial and frequency domain information across multiple scales to retain high-frequency components. We present a frequency-focused downsampling strategy to preserve critical spatial details during downsampling. Finally, we enhance the semantic representation capability of the model by aligning features from different feature fusion paths.
@@
    We propose UAV-DETR, an efficient end-to-end detector transformer for UAV imagery. The framework achieves superior accuracy and real-time performance. We put forward three models of different sizes to meet various performance requirements.
@@
    We develop a frequency-focused downsampling module that retains dual-domain information.
@@ III Methodology
As shown in Fig. 2, this study proposes a UAV-DETR, which is built upon the architecture of RT-DETR [6]. We enhance the model with three components, i.e., multi-scale feature fusion with frequency enhancement, frequency-focused downsampling, and semantic alignment and calibration. Additionally, we introduce inner Scylla intersection over union (Inner-SIoU) [18, 19] to replace the generalized intersection over union (GIoU).
Figure 2: Overview of the UAV-DETR. MSFF-FE represents the Multi-Scale Feature Fusion with Frequency Enhancement module; FD denotes Frequency-focused Downsampling; SAC is Semantic Alignment and Calibration. FFT and IFFT denote fast Fourier transform and its inverse operation, respectively.
@@ III-A Multi-Scale Feature Fusion with Frequency Enhancement
Figure 3: Structure of multi-scale feature fusion with frequency enhancement. By default, the number of output channels in the convolutional layers equals the number of input channels. The structural diagram of the Frequency-Focused module is included in Fig. 2.
In the proposed MSFF-FE architecture, the Focus module [20] is employed to compress spatial information from satellite imagery into channel dimensions through stride-based feature slicing and a convolution layer. As depicted in Fig. 3, we process S2 through Focus modules and concatenate the resulting feature map with other input feature maps along the channel dimension as input. The MSFF-FE adopts a cross-stage partial strategy [21]. It partitions the input feature map $x\in\mathbb{R}^{C\times H\times W}$ into two parts: $x_{1}\in\mathbb{R}^{C_{1}\times H\times W}$ and $x_{2}\in\mathbb{R}^{C_{2}\times H\times W}$. To avoid excessive computational cost, we set $C_{1}=\frac{C}{4}$ and $C_{2}=\frac{3C}{4}$. The first part $x_{1}$ undergoes multi-scale and frequency enhancement, while the second part $x_{2}$ is concatenated with the processed features from $x_{1}$, followed by a $1\times 1$ convolution to fuse the two branches. Firstly, the feature map $x_{1}$ passes through a $1\times 1$ convolution to adjust its channel dimensions, followed by the GELU activation function[22] for non-linearity, resulting in $x_{\text{conv}}$. As shown in Eq. (1), we transform $x_{\text{conv}}$ into the frequency domain by Fourier transform (F). To refine the frequency-domain features, we design a Global Average Pooling (GAP) layer, followed by a convolutional layer. After refinement, the features are transformed back to the spatial domain via inverse Fourier transform (IF).
@@
In our design, we also leverage a channel attention mechanism on $\text{GAP}(x_{\text{conv}})$ to further refine the $x_{\text{sc}}$. Then, We leverage the gating mechanism to modulate and refine the multi-scale features.
@@
where $\alpha$ and $\beta$ are learned parameters balancing the spatial and frequency components. We refer to this equation as the Frequency-Focused (FF) module, which will be used in later stages of the network. The FF module dynamically learns and adjusts filtering coefficients in the frequency domain through $1\times 1$ convolution layers. Additionally, we incorporate residual connections. This design preserves the original spatial features. At the same time, it allows for controllable fusion of frequency-enhanced components. Before combining with the unprocessed $x_{2}$, the enhanced features undergo final fusion.
@@
We employ large convolution kernels to capture long-range dependencies, while small convolution kernels are utilized to optimize channel-wise information. Additionally, residual connections are incorporated to accelerate model training. The final output is obtained by concatenating $x_{\text{final}}$ with $x_{2}$, followed by a $1\times 1$ convolution and GELU activation function.
III-B Frequency-Focused Downsampling
As illustrated in Fig. 2, in the frequency-focused downsampling (FD) module, the input feature map $x\in\mathbb{R}^{C\times H\times W}$ is first downsampled using average pooling with a kernel size of 2 and a stride of 1. As a result, the pooled feature map $x_{p}$ is obtained and then it is divided into two parts along the channel dimension: $x_{1}$ and $x_{2}$. Each of them are processed in parallel.
$x_{1}$ is processed using a $3\times 3$ convolution with stride 2 and padding 1, which reduces its spatial dimensions while preserving key features. Then, $x_{1}^{\prime}$ is obtained. $x_{2}$ undergoes parallel processing, where one path applies FF module to enhance important feature components. Accordingly, $x_{f}$ is obtained. The other path applies max pooling with a $3\times 3$ kernel and stride 2, followed by a $1\times 1$ convolution to reduce the number of channels. Correspondingly, we get $x_{p}^{\prime}$. The two outputs, $x_{f}$ and $x_{p}^{\prime}$, are concatenated along the channel dimension and passed through a $1\times 1$ convolution to reduce the number of channels to the desired size, which results in $x_{2}^{\prime}$. Finally, $x_{1}^{\prime}$ and $x_{2}^{\prime}$ are concatenated to form the final output of the module.
@@ III-C Semantic Alignment and Calibration
Given two input features $x_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ and $x_{2}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$, the SAC module first unifies the number of channels to a common dimension $C$ through separate convolution layers. Then, the feature $x_{2}$ is upsampled using bilinear interpolation to match the spatial dimensions of $x_{1}$. To enhance feature $x_{2}$, we apply FF module that generates a frequency-enhanced feature $x_{\text{freq}}$. Afterward, we fuse the frequency-enhanced feature with the original upsampled feature $x_{2}$. A gating mechanism is employed to balance the contributions from both the spatial and frequency domains.
Figure 4: Structure of semantic alignment and calibration.
@@
To address misalignment between $x_{1}$ and $x_{\text{fused}}$, the SAC module learns 2D offsets $\Delta_{1}$ and $\Delta_{2}$, which adjust the sampling grid for each feature map. These offsets are generated through a convolution layer. Using the learned offsets, we adjust the spatial coordinates of the features through a grid-based sampling operation [23] to ensure the alignment of both features.
@@
GridSample employs learnable offsets $\Delta$ to dynamically warp feature coordinates via bilinear interpolation to enable geometric alignment of multi-scale features. This differentiable formulation preserves gradient flow while adapting to viewpoint variations and scale discrepancies in UAV imagery.
Then, an element-wise weighted summation is used to fuse the aligned features, where $\alpha$ and $\beta$ are learned attention weights that balance the contributions from each aligned feature.
$x_{\text{output}}=\alpha\cdot x_{1}^{\text{aligned}}+\beta\cdot x_{\text{fused}}^{\text{aligned}}$ (8)
@@ III-D LOSS Function
RT-DETR uses GIoU loss for bounding box regression, which is less effective for small object detection, especially when the Intersection over Union (IoU) values are low. To address this issue, we adopt Inner-SIoU, which enhances both small object detection and geometric alignment. Inner-SIoU combines Inner-IoU[19] and SCYLLA-IoU (SIoU)[18]. It scales the auxiliary bounding box by 1.25 to improve sensitivity and speed up convergence. The InnerIoU is defined using expanded auxiliary bounding boxes. The Inner-IoU is calculated as:
@@ IV-A Experimental Setup
Datasets: We conduct quantitative experiments on two object detection datasets: VisDrone[24] and UAVVaste[25]. The VisDrone-2019-DET dataset comprises 6,471 training images, 548 validation images, and 3,190 test images, all captured from drones at varying altitudes across different locations. Each image is annotated with bounding boxes for ten predefined object categories: pedestrian, person, car, van, bus, truck, motorbike, bicycle, awning-tricycle, and tricycle. We use the VisDrone-2019-DET training set and validation set for training and testing, respectively.
@@
Implementation Details: All models are trained on NVIDIA GeForce RTX 3090. The UAV-DETR model is based on RT-DETR[6], with three variations: one using ResNet18, another with ResNet50, and a third incorporating EfficientFormerV2[26] as the backbone. We implement our approach and train the framework for 400 epochs with a batch size of 4. We use an early stopping mechanism with a patience setting of 20. The UAV-DETR network is optimized using AdamW[27] with a learning rate of 0.0001, a momentum of 0.9. We scale our input images to 640$\times$640 pixels and use the data augmentation methods from the RT-DETR model to ensure consistency across experiments. Additionally, we apply mixup [28] and Mosaic [29] techniques, with Mosaic set to a probability of 1 and mixup set to a probability of 0.2. We report the standard COCO metrics, including AP (averaged over uniformly sampled IoU thresholds ranging from 0.50-0.95 with a step size of 0.05), AP50, as well as AP at different scales: APS, APM.
@@
YOLOv9-S[30] ECCV2025 640×640 7.2 26.7 22.7 38.3
YOLOv9-M[30] ECCV2025 640×640 20.1 76.8 25.2 42.0
@@
HIC-YOLOv5[2] ICRA2024 640×640 9.4 31.2 26.0 44.3
@@
Model Params (M) GFLOPs APS APM AP AP50
YOLOv11-S[32] 9.4 21.3 27.3 48.7 27.8 63.0
HIC-YOLOv5[2] 9.4 31.2 30.8 20.9 30.5 65.1
RT-DETR-R18[6] 20.0 57.3 35.8 64.8 36.3 72.6
RT-DETR-R50[6] 42.0 129.9 37.0 62.3 37.4 73.5
UAV-DETR-EV2 (Ours) 13 43 36.7 63.3 37.1 70.6
UAV-DETR-R18 (Ours) 20 77 36.6 64.6 37.0 74.0
UAV-DETR-R50 (Ours) 42 170 37.1 61.3 37.5 75.9
IV-B Comparative Experiments
As listed in Table I, on the VisDrone dataset, our UAV-DETR-R18 achieves a 3.1% improvement in AP, a 4.2% increase in $\text{AP}_{50}$ compared to the baseline RT-DETR-R18. Similarly, the UAV-DETR-R50 sees a 3.1% increase in AP, a 4.1% rise in $\text{AP}_{50}$ compared to the baseline. UAV-DETR-EV2 also demonstrates an accuracy advantage over HIC-YOLOv5, which has a similar computational cost. UAV-DETR-R18 outperforms all methods with a computational cost below 100 GFLOPs, which achieves the best accuracy in its class. Furthermore, we compared our method with other object detectors that have computational costs similar to UAV-DETR-R50, and the results show that our approach also outperforms the others in terms of accuracy. Remarkably, our method still exhibits outstanding performance, even compared to approaches like PP-YOLOE-P2-Alpha-l [33], which typically benefit from extensive pre-training. Such comparisons are often considered unfair due to the significant advantages conferred by extensive pre-training. To further demonstrate the generalizability of UAV-DETR, we also evaluated the method on the UAVVaste dataset. The results are given in Table II. Notably, UAV-DETR still maintains a competitive advantage compared to other models. Compared to VisDrone, the UAVVaste dataset has a smaller amount of data. The strong performance of our model demonstrates that it does not rely on large amounts of annotated data.
@@ IV-B Comparative Experiments
RT-DETR-R50[6] 42 130 89
RT-DETR-R18[6] 20 60 183
UAV-DETR-R50 42 170 65
UAV-DETR-R18 20 77 124
UAV-DETR-EV2 13 43 116
IV-C Ablation Studies
We conducted ablation experiments on the VisDrone dataset using UAV-DETR-R18 to analyze the impact of each component on detection accuracy. Table III presents the performance comparison with different configurations, where IS stands for Inner-SIoU, MSFF-FE represents Multi-Scale Feature Fusion with Frequency Enhancement module, FD stands for Frequency-Focused Downsampling module, and SAC refers to Semantic Alignment and Calibration module.
The baseline RT-DETR-R18 achieves an AP of 26.7 and AP50 of 44.6. After incorporating Inner-SIoU, AP increased to 27.1, which shows that improving the loss function positively impacts performance. Adding the MSFF-FE module further improves the AP to 28.4, which demonstrates the benefits of incorporating multi-scale feature fusion and frequency enhancement. The addition of FD improves AP50 to 47.1. When all components are combined, UAV-DETR-R18 achieves the highest performance with an AP of 29.8 and an AP50 of 48.8, which showcases the cumulative impact of each module on detection accuracy. As given in Table IV, our experiments demonstrate that setting the ratio of Inner-SIOU to 1.25 is an appropriate choice. Lastly, we calculated the Frames Per Second (FPS) for both RT-DETR models and UAV-DETR models using the PyTorch implementation in 32-bit Floating Point Precision, as shown in Table V. The result demonstrates that UAV-DETR largely maintains the real-time performance of RT-DETR.